# New Multi-messenger Probe of Dark Matter-Nucleon Interactions from Ultra-high Energy Cosmic Ray Acceleration

### v1

@@ Abstract
It has been suggested that the density of dark matter (DM) halo can be highly enhanced around supermassive black holes at the centers of massive galaxies. If real, these DM spikes would offer new opportunities to probe the properties of DM. In this work, we point out that DM spikes can significantly impact the composition and survivability of ultra-high-energy cosmic rays accelerated near supermassive black holes. A large DM-nucleon cross section would fragment heavy nuclei into lighter elements and prevent them from attaining the energies observed at Earth. While the origin of cosmic rays remains a mystery, we show that if the highest-energy cosmic rays on Earth come from sources like NGC1068, then cross sections of size $\sigma_{\chi p}\leq 3\times 10^{-34}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}$ would be excluded by cosmic ray data. These bounds can be competitive with other existing probes in the DM mass region $m_{\chi}\in[5\;\mathrm{MeV},50\;\mathrm{MeV}]$. While the uncertainties on the acceleration mechanism of cosmic rays prevent us from setting robust limits, our study highlights an important connection between DM spikes and cosmic ray physics that is complementary to existing cosmological and direct detection constraints.
@@ I Introduction
Supermassive black holes (SMBH), believed to exist at the centers of almost all massive galaxies, provide extreme and naturally-occurring laboratories with which to probe the microphysics of cosmic rays (CRs), neutrinos, and dark matter (DM). These objects, with masses ranging from millions to billions of solar masses, create some of the most energetic environments in the Universe, where particles can be accelerated to ultra-high energies through various mechanisms including magnetic reconnection, shock acceleration, and interactions with accretion disk turbulence [Ptitsyna:2015nta, Inoue:2019yfs, Murase:2019vdl, Katsoulakos:2020rkr, Kheirandish:2021wkm, Eichmann:2022lxh, Inoue:2022yak, Fiorillo:2024akm]. The recent discovery of NGC 1068 [IceCube:2022der] as a high-energy neutrino source and its lack of corresponding TeV gamma-ray signal have shown that the close vicinity of the central SMBH is the most likely source of high-energy neutrino production [Padovani:2024ibi]. The suppression of gamma rays, combined with the fact that this is the first steady source of neutrino observed, suggests that the emission region is deeply embedded within dense material that effectively absorbs high-energy electromagnetic radiation while allowing neutrinos to escape. This, in turn, implies that CRs are accelerated near the central SMBH, within a few to hundreds of Schwarzschild radii from the event horizon. These small distances would imply that the CR acceleration is taking place in an environment with a potentially large ambient DM density, assuming a cuspy profile.
@@
Figure 1: The approximate bounds set here (pink region) using a spiked DM-density profile [Gondolo:1999ef] in an NGC 1068-type galaxy compared to other proposed bounds using CRs (black lines). Specifically, we are showing scattering [Cappiello:2018hsu], and cooling [Herrera:2023nww] bounds. We also show DM-proton cross section values that can produce the correct relic abundance via freeze-in (green region) [Elor:2021swj, Bhattiprolu:2022sdd]. Our potential bounds using a Burkert [Burkert:1995yz], NFW [Navarro:1995iw], and spiked-density profiles for Milky-Way-like galaxies are shown as dotted, dashed-dotted, and dashed lines, respectively.
@@
Gamma-ray observations from other galaxies (e.g., NGC 4486) have also demonstrated that high-energy emission can originate from regions very close to the event horizon, supporting the plausibility of particle acceleration at these scales [Berge:2006uls, Aharonian:2007ig, Acciari:2008ah, HESS:2011huh, MAGIC:2012gcb, Aleksic:2014xsg]. Such obscured, Compton-thick sources have long been theorized [Berezinsky:1977, Eichler:1979yy, Silberberg:1979hd, Hickox:2018xjf], but their existence is only now being confirmed with the advent of multi-messenger astronomy [IceCube:2021pgw, Giommi:2021bar, Fang:2022trf]. For instance, these obscured Active Galactic Nuclei (AGNs) have been used to explain the X-ray background (XRB) measurements [setti1989active], specifically the peak between 20-30 keV [Comastri:1994bz, Treister:2006ec, Akylas:2012am, Ananna:2018uec]. Interestingly, non-jetted AGNs have also been used to explain the XRB [Gilli:2006zi] and astrophysical neutrino background [Padovani:2024tgx]. Similar arguments have also been recently applied to NGC 7469 [Yang:2025lmb].
@@
While cuspy profiles are often used to describe DM halos, the presence of a SMBH can further enhance the DM density in its vicinity. It has been suggested that the DM density profile could form a ‘spike’ as we go closer to the SMBH due to adiabatic growth of the black hole [Gondolo:1999ef]. In other words, as the central black hole forms and grows, its gravity pulls in ambient DM, compressing it into a much steeper density profile, referred to as a “spike”. The origin of a spiked profile can be theoretically motivated simply from angular momentum conservation. With more mass accreting onto the SMBH, the minimum of the effective gravitational potential shifts towards the center, thus giving rise to a spike profile.
@@
Qualitatively, the spike formation can be understood as follows: Starting from an initial DM density profile $\rho(r)\propto r^{-\gamma}$ with $\gamma$ an initial density slope (typically $0.5\lesssim\gamma\lesssim 1.5$ [Gnedin:2003rj, 10.1093/mnras/stv058, 2017PDU....15...90I, Hooper:2016ggc, Baumgart:2025dov]), the adiabatic growth of the black hole causes the orbits of DM particles to contract towards the center. This process, which conserves adiabatic invariants, results in a new power-law density profile within the spike, $\rho^{\gamma_{\rm sp}}_{\rm sp}(r)\propto r^{-\gamma_{\rm sp}}$. The relationship between the initial slope $\gamma$ and the spiked slope $\gamma_{\rm sp}$ is given by $\gamma_{\rm sp}=\frac{9-\gamma}{4-\gamma}$ [Gondolo:1999ef, Quinlan:1994ed]. For typical halo models where $\gamma$ ranges from 0 to 2, this transformation yields very steep inner profiles, with $\gamma_{\rm sp}$ values between 2.25 and 2.5. The spike is theorized to extend from a minimum radius, influenced by annihilation or capture effects, out to a characteristic radius $R_{\rm sp}\sim 1\;\mathrm{pc}$. These results were then improved by using a fully relativistic treatment [Sadeghian:2013laa].
The potential existence of these spikes has been leveraged in numerous analyses to enhance expected sensitivities for DM searches, including studies by FGST and PAMELA [Sandick:2010yd], various gamma-ray observations [Regis:2008ij, Balaji:2023hmy], and stellar observations [Lacroix:2018zmg, Gustafson:2025ypo, Acevedo:2025rqu]. Furthermore, the spike has been utilized to set bounds on DM interactions through its capture and annihilation within stars [John:2023knt] and to predict the existence of dark stars [BetancourtKamenetskaia:2025hgr]. More recently [Akita:2025dhg] used the spike to explain the neutrino signal from NGC1068 via DM annihilation. Additionally, [Cline:2022qld, Ferrer:2022kei, Cline:2023tkp, Dev:2025tdv] have used spike profiles to constraint neutrino-DM interactions from IceCube and future galactic supernova neutrino observations. Collectively, these studies highlight the broad phenomenological relevance of spiked profiles. Note, however, that the existence of the spike is not yet established observationally and is still under debate [Ullio:2001fb, Lacroix:2018zmg, Shen:2023kkm]. See Appendix A for a review of possible mechanisms that could weaken or destroy such spikes.
@@
In this paper, we discuss how the existence of a spiked DM halo impacts CR acceleration to ultra-high energies. These ultra-high energy CRs (UHECRs) are expected predominantly to consist of heavier elements, including iron [Gaisser:2011klf, Thoudam:2016syr, Muzio:2019leu, Ivanov:2020rqn, PierreAuger:2023bfx, Ehlert:2023btz, Zhang:2024sjp, TelescopeArray:2024oux, IceCubeCollaborationSS:2025jbi, PierreAuger:2025eun, IceCube:2025baz]. A single scattering event between such energetic CRs and a DM particle during the acceleration process would cause the heavy (iron) nucleus to shatter. If such fragmentation occurs, CRs would be unable to reach the observed ultra-high energies, particularly near the Greisen-Zatsepin-Kuzmin (GZK) cutoff [Greisen:1966jv, Zatsepin:1966jv]. In fact, the observation of several super-GZK events [HIRES:1994ijd, PierreAuger:2022qcg, TelescopeArray:2023sbd] strongly suggest that (i) these UHECRs are most likely heavy nuclei (with higher GZK cutoff) rather than protons, and (ii) fragmentation of these heavy nuclei is unlikely to happen at the source.
@@
Here we ask the question: Can high-energy CRs be accelerated within the DM spike to the ultra-high energies necessary to explain current observations by CR experiments [Ivanov:2020rqn, PierreAuger:2020qqz]? We split this question into three parts:
@@
Given the observation of UHECRs on Earth, we can set bounds on the DM-nucleon cross section by requiring that the interaction length of CR-DM interactions is at least larger than the acceleration region. Figure 1 shows the approximate results we predict compared to other CR bounds. Not shown are bounds due to changes in the expected metallicity [Lu:2023aar] measured by CR experiments such as AMS [AMS:2021lxc]. There, it is assumed the DM-nucleon cross section behaves as the proton-nucleon cross section, scaled by a phenomenological factor $b_{\chi}$, $\sigma_{\chi-N}=b_{\chi}\sigma_{p-N}$. For instance, in some scenarios, the DM-nucleon cross section scales as $\sigma_{\chi-N}\sim A^{2/3}\ln^{2}(s)$, where $A$ is the nucleon count, and $\sqrt{s}$ the center of mass energy [Froissart:1961ux, Martin:1962rt, Glauber:1970jm, Block:2005pt, Ulrich:2010rg, Gaisser:2013bla]. With such a scaling, we would have to improve our bounds by $\mathcal{O}(100)$ to compare with [Lu:2023aar] and $\mathcal{O}(1000)$ to compare with direct detection experiments. Since we wish to remain agnostic to the underlying model, we assume $\sigma_{\chi-N}$ is constant with the momentum transfer.
Note, that the method presented here is similar to [Cappiello:2018hsu, Lu:2023aar]. The main difference is that the authors focused on changes in the energy spectrum (at lower energies) due to multiple scattering. In contrast, here we focus on single-scatter events at higher energies.
This article is structured as follows. In Section II, we briefly discuss the DM-density profiles used in this work and in Section III we discuss general CR-DM interactions. Then, in Section IV, we derive bounds using EeV CR iron nuclei by requiring that it does not fragment during acceleration to these energies. We present our conclusions in Section V.
@@ II DM Densities
Figure 2: The DM density as predicted by [Gondolo:1999ef] combined with the NFW [Navarro:1995iw] profile starting at $r=4R_{\mathrm{S}}=1.6\times 10^{-9}\;\mathrm{kpc}$. The orange region shows where DM self-interactions (e.g., annihilation) could reduce the expected density. The black dashed line denotes the expected DM density when accounting for significant stellar heating [Balaji:2023hmy], and the dotted black line indicates the plateau induced by an annihilation of $\langle\sigma v\rangle\sim 10^{-25}$ cm3/s. The pink vertical lines indicate the conservative acceleration regions of CRs for a Milky-Way-like or NGC1068-like galaxy, respectively, based on multi-messenger observations.
To model the DM density profile near the SMBH, we follow the formalism of [Gondolo:1999ef]. We adopt a generalized NFW profile [Navarro:1995iw] for the ambient dark-matter halo outside of the spike region, with an initial slope parameter $\gamma$ set to 1 and $\rho_{0}=0.4\;\mathrm{GeV/cm^{3}}$, the local ($r_{0}=8.5\;\mathrm{kpc}$) dark matter density. In the simplest scenario, where the central SMBH grows adiabatically, via the slow in-fall of gas, the DM density can be written as
@@
Here $R_{\rm S}=2GM_{\rm BH}$ is the Schwarzschild radius of the black hole, $R_{\rm sp}=\alpha_{\gamma}r_{0}(M_{\rm{BH}}/(\rho_{0}r^{3}))^{1/(3-\gamma)}$ the size of the spike, and $\alpha_{\gamma}$ the normalization to match the density outside of the spike (here for $\gamma=1\rightarrow\alpha_{\gamma}=0.12,\gamma_{\rm{sp}}=7/3$). Thus, the full density profile is
@@
where $\rho^{7/2}_{\rm sp}$ is the DM density with the spike parameter $\gamma_{\rm sp}=7/2$. There are multiple ways the above adiabatic accumulation of DM could be perturbed. For example, first, stars orbiting the central black hole can soften the spike via gravitational heating of the DM, and, if significant, can change the expected DM profile. This is reflected by the change in expected $\gamma_{\rm sp}$. For the Milky-Way with the profile given above $\gamma_{\rm sp}=7/3\rightarrow\gamma_{\rm sp}=3/2$ [Bertone:2005hw, Bertone:2005xv, Balaji:2023hmy]. Secondly, if DM self-interacts, e.g. via annihilation, then the maximum of the spike would be reduced to $\rho_{c}$ [Gondolo:1999ef, Balaji:2023hmy, Akita:2025dhg]. Here $\rho_{\rm c}=m_{\chi}/(\langle\sigma v\rangle t_{\rm BH})$ with $m_{\chi}$ the DM mass, $\langle\sigma v\rangle$ the thermally averaged annihilation cross-section, and $t_{\rm BH}$ the age of the SMBH. Combined, the resulting DM distribution (for the Milky-Way) would be
@@
where $\rho^{3/2}_{\rm sp}$ stands for the DM density in the spike with $\gamma_{\rm sp}=3/2$. Note, while in the above equations we set the lower spike bound to be $4R_{\mathrm{S}}$, the lower bound could be as small as $2R_{\mathrm{S}}$ [Sadeghian:2013laa, Ferrer:2017xwm, Ferrer:2022kei].
Figure 2 shows the DM density profiles for the MW used in this work. We set $M_{\rm BH}=4\times 10^{6}~M_{\odot}$, and $\gamma=1$. Above one pc, the profile follows a standard NFW profile (green dashed), whereas below one pc, the spike takes over (solid black). The dotted line denotes the effects of self-annihiltion when setting $\langle\sigma v\rangle=10^{-25}\;\mathrm{cm^{3}/s}$, while the black dashed line denotes the density profile when including stellar heating.
In the following we discuss CR acceleration within Milky-Way-like galaxies ($M_{\mathrm{BH}}\sim 4\times 10^{6}M_{\odot}$ [GRAVITY:2023avo]), and NGC 1068-like galaxies ($M_{\mathrm{BH}}\sim 8\times 10^{6}M_{\odot}$ [Lodato:2002cv]). This puts them at the low end of expected masses for SMBHs in AGNs, where we expected ranges from $10^{6}M_{\odot}-10^{10}M_{\odot}$ [Woo:2002un, Dalla_Bont__2025]. Since the spike density scales with the mass of the SMBH as $\sim M^{0.7}$, using low masses, as we do here, should be considered a conservative approach.

### v2

@@ Abstract
It has been suggested that the density of dark matter (DM) halo can be highly enhanced around supermassive black holes at the centers of massive galaxies. If real, these DM spikes would offer new opportunities to probe the properties of DM. In this work, we point out that DM spikes can significantly impact the composition and survivability of ultra-high-energy cosmic rays accelerated near supermassive black holes. A large DM-nucleon cross section would fragment heavy nuclei into lighter elements and prevent them from attaining the energies observed at Earth. While the origin of cosmic rays remains a mystery, we show that if the highest-energy cosmic rays on Earth come from sources like NGC 1068, then cross sections of size $\sigma_{\chi p}\leq 10^{-33}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}$ would be excluded by cosmic ray data. These bounds can be competitive with other existing probes and rule out new parameter space in the DM mass region $m_{\chi}\in[3\;\mathrm{MeV},30\;\mathrm{MeV}]$. While the uncertainties on the acceleration mechanism of cosmic rays prevent us from setting robust limits, our study highlights an important connection between DM spikes and cosmic ray physics that is complementary to existing cosmological and direct detection constraints.
@@ I Introduction
Supermassive black holes (SMBH), believed to exist at the centers of almost all massive galaxies, provide extreme and naturally-occurring laboratories with which to probe the microphysics of cosmic rays (CRs), neutrinos, and dark matter (DM). These objects, with masses ranging from millions to billions of solar masses, create some of the most energetic environments in the Universe, where particles can be accelerated to ultra-high energies through various mechanisms including magnetic reconnection, shock acceleration, and interactions with accretion disk turbulence [142, 107, 131, 115, 116, 75, 106, 85, 84, 114]. The recent discovery of NGC 1068 [4] as a high-energy neutrino source and its lack of corresponding TeV gamma-ray signal have shown that the close vicinity of the central SMBH is the most likely source of high-energy neutrino production [138]. The suppression of gamma rays, combined with the fact that this is the first steady source of neutrino observed, suggests that the emission region is deeply embedded within dense material that effectively absorbs high-energy electromagnetic radiation while allowing neutrinos to escape. This, in turn, implies that CRs are accelerated near the central SMBH, within a few to hundreds of Schwarzschild radii from the event horizon. These small distances would imply that the CR acceleration is taking place in an environment with a potentially large ambient DM density, assuming a cuspy profile.
@@
Figure 1: The approximate bounds set here (pink region) using a spiked DM-density profile [96] in an NGC 1068-type galaxy compared to other proposed bounds using CRs (black lines). Specifically, we are showing scattering [58], and cooling [102] bounds. We also show DM-proton cross section values that can produce the correct relic abundance via freeze-in (green region) [76, 52]. Our potential bounds using a Burkert [57], NFW [134], and spiked-density profiles for Milky-Way (MW)-like galaxies are shown as dotted, dashed-dotted, and dashed lines, respectively.
@@
Gamma-ray observations from other galaxies (e.g., NGC 4486) have also demonstrated that high-energy emission can originate from regions very close to the event horizon, supporting the plausibility of particle acceleration at these scales [45, 23, 16, 14, 26, 27]. Such obscured, Compton-thick sources have long been theorized [44, 74, 158, 103], but their existence is only now being confirmed with the advent of multi-messenger astronomy [5, 92, 81]. For instance, these obscured Active Galactic Nuclei (AGNs) have been used to explain the X-ray background (XRB) measurements [155], specifically the peak between 20-30 keV [64, 162, 25, 28]. Interestingly, non-jetted AGNs have also been used to explain the XRB [91] and astrophysical neutrino background [136]. Similar arguments have also been recently applied to NGC 7469 [170].
@@
While cuspy profiles are often used to describe DM halos, the presence of an SMBH can further enhance the DM density in its vicinity. It has been suggested that the DM density profile could form a ‘spike’ as we go closer to the SMBH due to adiabatic growth of the black hole [96]. In other words, as the central black hole forms and grows, its gravity pulls in ambient DM, compressing it into a much steeper density profile, referred to as a “spike”. The origin of a spiked profile can be theoretically motivated simply from angular momentum conservation. With more mass accreting onto the SMBH, the minimum of the effective gravitational potential shifts towards the center, thus giving rise to a spiked profile.
@@
Qualitatively, the spike formation can be understood as follows: Starting from an initial DM density profile $\rho(r)\propto r^{-\gamma}$ with $\gamma$ an initial density slope (typically $0.5\lesssim\gamma\lesssim 1.5$ [95, 141, 108, 105, 42]), the adiabatic growth of the black hole causes the orbits of DM particles to contract towards the center. This process, which conserves adiabatic invariants, results in a new power-law density profile within the spike, $\rho^{\gamma_{\rm sp}}_{\rm sp}(r)\propto r^{-\gamma_{\rm sp}}$. The relationship between the initial slope $\gamma$ and the spiked slope $\gamma_{\rm sp}$ is given by $\gamma_{\rm sp}=\frac{9-\gamma}{4-\gamma}$ [96, 144]. For typical halo models where $\gamma$ ranges from 0 to 2, this transformation yields very steep inner profiles, with $\gamma_{\rm sp}$ values between 2.25 and 2.5. The spike is theorized to extend from a minimum radius, influenced by annihilation or capture effects, out to a characteristic radius $R_{\rm sp}\sim 1\;\mathrm{pc}$. These results were then improved by using a fully relativistic treatment [151]. The potential existence of these spikes has been leveraged in numerous analyses to enhance expected sensitivities for DM searches, including studies by FGST and PAMELA [152], various gamma-ray observations [145, 97, 40], and stellar observations [120, 101, 17]. Furthermore, the spike has been utilized to set bounds on DM interactions through its capture and annihilation within stars [112] and to predict the existence of dark stars [51]. Refs. [166, 98] have used spikes around the TXS 0506+056 to constrain blazar-boosted DM. More recently, Ref. [24] used the spike to explain the neutrino signal from NGC 1068 via DM annihilation. Additionally, Refs. [62, 83, 63, 69, 172, 70, 72, 68] have used spiked profiles to constraint neutrino-DM interactions. Collectively, these studies highlight the broad phenomenological relevance of spiked profiles. Note, however, that the existence of the spike is not yet established observationally and is still under debate [164, 120, 156]. See Appendix A for a review of possible mechanisms that could weaken or destroy such spikes.
@@
In this paper, we discuss how the existence of a spiked DM halo impacts CR acceleration to ultra-high energies. These ultra-high energy CRs (UHECRs) are expected predominantly to consist of heavier elements, including iron [89, 161, 132, 109, 126, 73, 174, 9, 7, 13, 6]. A single scattering event between such energetic CRs and a DM particle during the acceleration process would cause the heavy (iron) nucleus to shatter. If such fragmentation occurs, CRs would be unable to reach the observed ultra-high energies, particularly near the Greisen-Zatsepin-Kuzmin (GZK) cutoff [99, 173]. In fact, the observation of several super-GZK events [54, 12, 8] strongly suggests that (i) these UHECRs are most likely heavy nuclei (with higher GZK cutoff) rather than protons, and (ii) fragmentation of these heavy nuclei is unlikely to happen at the source.
@@
Here we ask the question: Given a DM-proton interaction, can high-energy CRs be accelerated within the DM spike to the ultra-high energies necessary to explain current observations by CR experiments [109, 1]? We split this question into three parts:
@@
Given the observation of UHECRs on Earth, we can set bounds on the DM-nucleon cross section by requiring that the interaction length of CR-DM interactions is at least larger than the acceleration region. Figure 1 shows the approximate results we predict compared to other CR bounds. Not shown are bounds due to changes in the expected metallicity [124] measured by CR experiments such as AMS [20]. There, it is assumed the DM-nucleon cross section behaves as the proton-nucleon cross section, scaled by a phenomenological factor $b_{\chi}$, $\sigma_{\chi N}=b_{\chi}\sigma_{pN}$. For instance, in some scenarios, the DM-nucleon cross section scales as $\sigma_{\chi N}\sim A^{2/3}\ln^{2}(s/{\rm GeV}^{2})$, where $A$ is the nucleon count, and $\sqrt{s}$ the center-of-mass energy [87, 125, 93, 55, 165, 88]. With such a scaling, our bounds would improve by $\mathcal{O}(100)$ when compared with Ref. [124] and by $\mathcal{O}(1000)$ when compared with direct detection experiments. Since we wish to remain agnostic to the underlying model, we assume $\sigma_{\chi N}$ is constant with the momentum transfer.
Note that the method presented here is similar to Refs. [58, 124]. The main difference is that those authors focused on changes in the energy spectrum (at lower energies) due to multiple scattering. In contrast, here we focus on single-scatter events at higher energies.
This article is structured as follows. In Section II, we briefly discuss the DM-density profiles used in this work. In Section III we discuss general CR-DM interactions. Then, in Section IV, we derive bounds on DM-proton interaction rate using EeV CR iron nuclei by requiring that they do not fragment during acceleration to these energies. We present our conclusions in Section V.
@@ II DM Densities
Figure 2: The DM density as predicted by Ref. [96] combined with the NFW [134] profile starting at $r=4R_{\mathrm{S}}=1.6\times 10^{-9}\;\mathrm{kpc}$ for an SMBH mass of $4\times 10^{6}M_{\odot}$. The orange region shows where DM annihilation could reduce the expected density. The black dashed line denotes the expected DM density when accounting for significant stellar heating [40], and the dotted black line indicates the plateau induced by DM annihilation of $\langle\sigma v\rangle\sim 10^{-25}$ cm3/s. While the density profile shown is for an MW-like galaxy, the pink vertical lines indicate the conservative acceleration regions of CRs for an MW-like or NGC 1068-like galaxy, respectively, based on multi-messenger observations.
To model the DM density profile near the SMBH, we follow the formalism of Ref. [96]. We adopt a generalized NFW profile [134] for the ambient DM halo outside of the spike region, with an initial slope parameter $\gamma$ set to 1 and $\rho_{0}=0.4\;\mathrm{GeV/cm^{3}}$, the local ($r_{0}=8.5\;\mathrm{kpc}$) DM density. In the simplest scenario, where the central SMBH grows adiabatically, via the slow in-fall of gas, the DM density can be written as [96]
@@
Here $R_{\rm S}=2GM_{\rm BH}$ is the Schwarzschild radius of the black hole, $R_{\rm sp}=\alpha_{\gamma}r_{0}(M_{\rm{BH}}/(\rho_{0}r^{3}))^{1/(3-\gamma)}$ the size of the spike, and $\alpha_{\gamma}$ the normalization to match the density outside of the spike. Here $\gamma=1$ corresponds to $\alpha_{\gamma}=0.12,\gamma_{\rm{sp}}=7/3$. Thus, the full density profile is
@@
where $\rho^{7/2}_{\rm sp}$ is the DM density with the spike parameter $\gamma_{\rm sp}=7/2$. There are multiple ways the above adiabatic accumulation of DM could be perturbed. For example, first, stars orbiting the central black hole can soften the spike via gravitational heating of the DM, and, if significant, can change the expected DM profile. This is reflected by the change in expected $\gamma_{\rm sp}$. For the MW with the profile given above, $\gamma_{\rm sp}=7/3\rightarrow\gamma_{\rm sp}=3/2$ [48, 47, 40, 50]. Secondly, if DM self-interacts, e.g. via annihilation, then the maximum of the spike would be reduced to $\rho_{c}$ [96, 40, 24]. Here $\rho_{\rm c}=m_{\chi}/(\langle\sigma v\rangle t_{\rm BH})$ with $m_{\chi}$ the DM mass, $\langle\sigma v\rangle$ the thermally-averaged annihilation cross-section, and $t_{\rm BH}$ the age of the SMBH. Combined, the resulting DM distribution for the MW would be
@@
where $\rho^{3/2}_{\rm sp}$ stands for the DM density in the spike with $\gamma_{\rm sp}=3/2$. Note, while in the above equations we set the onset of the spike starting at $4R_{\mathrm{S}}$, the lower bound could be as small as $2R_{\mathrm{S}}$ after taking into account general-relativistic effects [151, 82, 83], but it will have a negligible impact on our results.
Figure 2 shows the DM density profiles for the MW used in this work. We set $M_{\rm BH}=4\times 10^{6}~M_{\odot}$, and $\gamma=1$. Above one pc, the profile follows a standard NFW profile (green dashed), whereas below one pc, the spike takes over (solid black). The dotted line denotes the effects of self-annihiltion when setting $\langle\sigma v\rangle=10^{-25}\;\mathrm{cm^{3}/s}$, while the black dashed line denotes the density profile when including stellar heating [40].
In the following, we discuss CR acceleration within MW-like galaxies ($M_{\mathrm{BH}}\sim 4\times 10^{6}M_{\odot}$ [15]), and NGC 1068-like galaxies ($M_{\mathrm{BH}}\sim 8\times 10^{6}M_{\odot}$ [123]). This puts them at the low end of expected masses for SMBHs in AGNs, where we expected ranges from $10^{6}M_{\odot}-10^{10}M_{\odot}$ [167, 65]. Since the spike density scales with the mass of the SMBH as $\sim M^{0.7}$, using low masses, as we do here, should be considered a conservative approach. The corresponding acceleration regions are shown by the pink vertical lines in Figure 2.